OpenStudy (anonymous):

Find the x-intercept and y-intercept of the line given by the equation 4x-7y =9

OpenStudy (anonymous):

set x = 0 and solve for y and y = 0 and solve for x

OpenStudy (anonymous):

y = -9/7 and x = 9/4?

OpenStudy (anonymous):

yeah so what are the intercepts?